A load W = 666 Kilonewtons is supported by two steel cables one of which makes an angle of 35 with the horizontal while the other makes an angle of 75

with the horizontal.


What is the Force or tension in NEWTONS of the 35 cable?


1
Expert's answer
2021-03-21T11:31:20-0400

"T_1\cos35°=T_2\cos75°"


"T_1\sin35°+T_2\sin75°=W"


"T_2=T_1\frac{\cos35°}{\cos75°}=3.165\cdot T_1"


"T_1\sin35°+3.165\cdot T_1\cdot\sin75°=666000\to"


"T_1=183434\ (N)\approx183.4\ (kN)" . Answer